# Advancing Sequential Numerical Prediction in Autoregressive Models

### v1

@@ Abstract
1   Jinghui Lu∗1   Qi Sun2   Hao Feng†1   Yanjie Wang1
@@ 2 Related Work
The Earth Mover’s Distance (EMD) measures the minimal cost of transforming one distribution into another, and has become a valuable metric in deep learning applications. Notably, Wasserstein GAN Arjovsky et al. (2017) uses EMD as its loss function to stabilize training in GANs. Cuturi (2013) and  Courty et al. (2016) also adopted EMD for smoothing the training procedure. Despite the success of EMD, it has not been applied to autoregressive models. Most recently, autoregressive models, especially LLMs, have advanced NLP (Radford et al., 2019; Touvron et al., 2023; Lu et al., 2023b), and multi-modal tasks (Alayrac et al., 2022; Wang et al., 2024c). While the tasks mentioned above require high precision in numerical value prediction, none of the previous works have specifically optimized for this criterion. Our work addresses this gap by focusing on advancing the sequential numerical prediction for autoregressive models.
@@ 5 Conclusion
The limitations of the NTIL include its degradation to EMD loss when predicting a single token, which diminishes its effectiveness for broader sequence-level tasks. Additionally, the exponential position-based weighting scheme, while effective in many cases, had limited or negative impact in certain configurations, such as with the Mathvision dataset and LLaVA-1.5 model. Future exploration could focus on refining the exponential position-based weighting scheme with adaptive strategies to address its inconsistent impact.

### v2

@@ Abstract
1   Jinghui Lu∗1   Qi Sun∗2   Hao Feng†1   Yanjie Wang1
@@ 2 Related Work
The Earth Mover’s Distance (EMD) measures the minimal cost of transforming one distribution into another, and has become a valuable metric in deep learning applications. Notably, Wasserstein GAN Arjovsky et al. (2017) uses EMD as its loss function to stabilize training in GANs. Cuturi (2013) and  Courty et al. (2016) also adopted EMD for smoothing the training procedure. Despite the success of EMD, it has not been applied to autoregressive models. Most recently, autoregressive models, especially LLMs, have advanced NLP (Radford et al., 2019; Touvron et al., 2023; Lu et al., 2023b; Cui et al., 2025), and multi-modal tasks (Alayrac et al., 2022; Wang et al., 2024c; Feng et al., 2025; Lu et al., 2025). While the tasks mentioned above require high precision in numerical value prediction, none of the previous works have specifically optimized for this criterion. Our work addresses this gap by focusing on advancing the sequential numerical prediction for autoregressive models.
@@ 5 Conclusion
The limitations of the NTIL include the exponential position-based weighting scheme, while effective in many cases, had limited or negative impact in certain configurations, such as with the Mathvision dataset and LLaVA-1.5 model. Future exploration could focus on refining the exponential position-based weighting scheme with adaptive strategies to address its inconsistent impact.
Furthermore, NTIL introduces additional computational overhead compared to CE loss, resulting in reduced training efficiency. This trade-off between performance improvement and computational cost needs to be considered in practical applications.
Another limitation arises from the tokenization strategies of certain models like LLaMA-3, which encode common multi-digit numbers (e.g., "123") as a single token. Such cases require special handling in NTIL’s implementation, adding complexity to the framework.
@@ D.3 Predicting Digits with EMD
[…]
0 0.780 0.783 0.789 0.789 0.771
0.1 0.777 0.790 0.799 0.789 0.778
0.2 0.782 0.793 0.795 0.785 0.784
0.3 0.788 0.783 0.786 0.785 0.781
0.4 0.774 0.782 0.778 0.781 0.784
Table 7: Ablation studies of hyper-parameters $\alpha$ and $\beta$
0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7
$\sigma$ 0.784 0.785 0.789 0.788 0.783 0.780 0.770 0.763
$\lambda$ 0.741 0.777 0.798 0.795 0.788 0.786 0.787 0.777
Table 8: Ablation studies of the hyper-parameters $\sigma$ (exp-weighting) and $\lambda$ (coefficient)
Model Params CE EMD NTIL
Qwen2.5 2b 0.509 0.517 0.523
MiniCPM3 4b 0.672 0.676 0.704
Yi 6b 0.375 0.383 0.400
LLaMA3 8b 0.638 0.643 0.646
Table 9: Results of the GSM8k mathematical reasoning task.
@@ Appendix E Ablations
We have supplemented our work with comprehensive experiments on the clock time recognition task using Qwen2-VL-7B, as shown in Table 7 and Table 8. We have bolded some of the outperforming results. The results demonstrate that NTIL adapts well to different hyperparameters.
We have also supplemented NTIL with the GSM8k math reasoning dataset. NTIL brings consistent performance improvement on mixed text-numeric tasks, which shows the effectiveness of our method.